OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1). A. y + 1 = -2(x – 5) B. y – 1 = -2(x – 5) C. y – 5 = -2(x + 1) D. y -5 = -2(x – 1)

OpenStudy (ivettef365):

the formula is (y - y1) = m(x - x1) m = slope, substitute the points what do you get ?

OpenStudy (anonymous):

y -5 = -2(x – 1) ?

OpenStudy (ivettef365):

you have y - (-1) = -2 (x -5) points (5,-1) (x,y)